AD=600000 + .2IT IT=550000+.5AD solve for IT and AD

AD=600000 + .2IT

IT=550000+.5AD

solve for IT and AD
Answer #1

First equation : AD = 600000 + 0.2 IT

Second equation: IT = 550000 + 0.5 AD

Putting the value of IT in the first equation, we get,

AD = 600000 + 0.2 * (550000 + 0.5 AD)

AD = 600000 + 110000 + 0.1 AD

AD = 710000 + 0.1AD

AD - 0.1AD = 710000

0.9 AD = 710000

AD = 710000 / 0.9

AD = 788888.89

Now, putting the above calculated value of AD in the second equation, we get,

IT = 550000 + (0.5 * 788888.89)

IT = 550000 + 3944444.44

IT = 944444.44